An individual firm supply curve for the quantity of Beer supplied in litter per- hour is given by Qs = 60P - 48. There are 5 firms in the Beers industry and the price of Beer is 40 Birr per-litter. Find; A) the quantity of Beer supplied by an individual firm per- hour. B) The quantity of Beer supplied by the industry per- hour


1
Expert's answer
2022-01-05T11:13:54-0500

Solution:

A.). The quantity of beer supplied by an individual firm per hour:

Qs = 60P – 48

Qs = 60 (40) – 48 = 2,400 – 48 = 2,352

The quantity of beer supplied by an individual firm per hour = 2,352 units.


 

B.). The quantity of beer supplied by the industry per hour:

Qs = (60P – 48) x 5 = 300P – 240

Industry supply curve: Qs = 300P – 240

Qs = 300 (40) – 240 = 12,000 – 240 = 11,760

Qs = 11,760 units